# TimeFormer: Capturing Temporal Relationships of Deformable 3D Gaussians for Robust Reconstruction

### v1

@@ 3 Preliminary: Deformable 3D Gaussians
The covariance matrix $\Sigma$ can be decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$: $\Sigma=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{S}=\text{diag}([s_{x},s_{y},s_{z}])$ and $\mathbf{R}$ can be transformed from a quaternion $[r_{w},r_{x},r_{y},r_{z}]$. Then the 3D Gaussians can be splatted to a 2D camera plane through differential Gaussian splatting. To model the appearance of 3D Gaussians, spherical harmonics (SH) are introduced to define the color $c$. Finally, for each pixel, the rendering results of 3DGS can be derived by calculating the color contribution of all the related Gaussians. This process is known as $\alpha$-blending:
@@ 4.2 TimeFormer
$\gamma(p)=$ $\left(\sin\left(2^{0}\pi p\right),\cos\left(2^{0}\pi p\right),\right.\cdots$ (4)
@@
In our experiments, we set $L=6$ for both positions $x$ and time $t$. We treat $[\gamma(\mathcal{G}_{c},\gamma(\mathcal{T^{\prime}})]\in\mathbb{R}^{B\times N\times(3\times 2L+2L)}$ as the original input $F_{0}$ to TimeFormer. It can be considered as $N$ sequences of the length $B$, containing $8L$ feature channels. With $M$ transformer encoder layers, for $m^{th}$ layer, intermediate features are calculated through multi-head self-attention(MSA) and MLP blocks. In the final stage, we use a tiny MLP to transform the last encoded features $F_{M-1}$ into offset $\mathcal{O}\in\mathbb{R}^{B\times N\times 3}$ in the linear space:
@@ 4.3 Shared Deformation Fields
Method $N$downarrow Training $\downarrow$ FPS $\uparrow$
@@
Method $N$downarrow Training $\downarrow$ FPS $\uparrow$
@@ 5.2 Experimental Comparisons
Analysis of Motion Patterns. We argue that TimeFormer achieves more robust learning of motion patterns. To take a deeper insight, we visualize the motion changes as follows: for each Gaussian, we calculate its bias $(\Delta_{x},\Delta_{y},\Delta_{z})$ towards canonical space in time $t$ and replace the color attributes $(r,g,b)$ with the absolute value of this bias: $(r,g,b)\leftarrow(\lvert\Delta_{x}\rvert,\lvert\Delta_{y}\rvert,\lvert\Delta_{z}\rvert)$. We use the same splatting algorithm to render the accumulated motion bias, transferring the rendered result as a heatmap, as in Fig. 10.
@@ Appendix A Implementation
The Transformer Encoder receives input structured as [seq_len, seq_batch, channel], and we input x structured as [$B$, $N$, 4], where $B$ is the size of time batch, $N$ is the number of Gaussians and 4 means 3 position channel and 1 channel for the timestamp. Alg. 2 shows how we construct input to time.
class TimeFormer (nn.module):
@@
layer = nn.TransformerEncoderLayer(PE_ch, nhead, d_hidden, activation=nn.functional.tanh)
self.encoder = nn.TransformerEncoder(layer, n_layer)

### v2

@@ 3 Preliminary: Deformable 3D Gaussians
The covariance matrix $\Sigma$ can be decomposed into a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$: $\Sigma=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{S}=\text{diag}([s_{x},s_{y},s_{z}])$ and $\mathbf{R}$ can be transformed from a quaternion $r_{w},r_{x},r_{y},r_{z}]$. Then the 3D Gaussians can be splatted to a 2D camera plane through differential Gaussian splatting. To model the appearance of 3D Gaussians, spherical harmonics (SH) are introduced to define the color $c$. Finally, for each pixel, the rendering results of 3DGS can be derived by calculating the color contribution of all the related Gaussians. This process is known as $\alpha$-blending:
@@ 4.2 TimeFormer
$\gamma(p)=$ $\left(\sin\left(2^{0}\pi p\right),\cos\left(2^{0}\pi p\right),\right$$\cdots$ (4)
@@
In our experiments, we set $L=6$ for both positions $x$ and time $t$. We treat $\gamma(\mathcal{G}_{c},\gamma(\mathcal{T^{\prime}})]\in\mathbb{R}^{B\times N\times(3\times 2L+2L)}$ as the original input $F_{0}$ to TimeFormer. It can be considered as $N$ sequences of the length $B$, containing $8L$ feature channels. With $M$ transformer encoder layers, for $m^{th}$ layer, intermediate features are calculated through multi-head self-attention(MSA) and MLP blocks. In the final stage, we use a tiny MLP to transform the last encoded features $F_{M-1}$ into offset $\mathcal{O}\in\mathbb{R}^{B\times N\times 3}$ in the linear space:
@@ 4.3 Shared Deformation Fields
Method NN $\downarrow$ Training $\downarrow$ FPS $\uparrow$
@@
Method NN $\downarrow$ Training $\downarrow$ FPS $\uparrow$
@@ 5.2 Experimental Comparisons
[…]
@@ Appendix A Implementation
The Transformer Encoder receives input structured as [seq_len, seq_batch, channel], and we input x structured as [$B$, NN, 4], where $B$ is the size of time batch, $N$ is the number of Gaussians and 4 means 3 position channel and 1 channel for the timestamp. Alg. 2 shows how we construct input to time.
class TimeFormer ($n$.module):
@@
layer = $n$.TransformerEncoderLayer(PE_ch, nhead, d_hidden, activation=$n$.functional.tanh)
self.encoder = $n$.TransformerEncoder(layer, n_layer)